OpenStudy (anonymous):

Find the solutions of the equation that are in the interval [0, 2π). 2-2sin(t)=2sqrt(3)cos(t)

OpenStudy (anonymous):

Consider squaring both sides and using Pythagorean identity.

OpenStudy (anonymous):

Should I factor out the 2 on the left and cancel first?

OpenStudy (anonymous):

yeah that could help.

OpenStudy (anonymous):

Ok, well, if I square the left I will get sin^2x-2sinx+1, and that isn't very helpful from what I can tell right?

OpenStudy (anonymous):

and then on the left have 3cos^2x

OpenStudy (anonymous):

sorry for mixing my variable name up

OpenStudy (anonymous):

\[ 3-3\cos^2(t)=3\sin^2(t) \]

OpenStudy (anonymous):

I'm sorry, I don't understand. 2(1-sint)=2sqrt(3)cost 1-sint=sqrt(3)cost Now if I square both sides (1-sint)^2 does not equal 1-sin^2t but 1-2sint+sin^2t

OpenStudy (anonymous):

I never said it did.

OpenStudy (anonymous):

How'd you get 3-3cos^2t=3sin^2t?

OpenStudy (anonymous):

It's an equation, you can add it whatever you want to add in if you do it to both sides.

OpenStudy (anonymous):

subtract \(3\) from both sides. \(3\cos^2(t)-3 = - (3-3\cos^2(t))\)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

This looks sketchy without seeing your exact steps

OpenStudy (anonymous):

\[\begin{split} 1-2\sin(t)+\sin^2(t)&=3\cos^2(t)\\ -2-2\sin(t)+\sin^2(t)&=3\cos^2(t)-3\\ -2-2\sin(t)+\sin^2(t)&=-(3-3\cos^2(t))\\ -2-2\sin(t)+\sin^2(t)&=-3\sin^2(t)\\ -2-2\sin(t)+4\sin^2(t)&=0&x=\sin(t)\\ 4x^2-2x-2&=0 \end{split}\]

OpenStudy (anonymous):

Ok! This makes a lot of sense now. I guess I was confusing myself too much by focusing so hard on the left hand side when the right hand side was easier.